OpenStudy (anonymous):

Using quadratics, solve: The floor of a shed has an area of 96 square feet. The floor is in the shape of a rectangle whose length is 4 feet less than twice the width. Find the length and the width of the floor of the shed.

OpenStudy (anonymous):

nm sloved it w=8 l =12

OpenStudy (anonymous):

A = w * ( 2w - 4) = 96 2w^2 - 4w - 96 = 0 w^2 - 2w - 48 = 0 ( w -8 ) ( w + 6) = 0 => w = 8ft , L = 12 ft